calculate the speed of sound in helium gas at STP.
1
Expert's answer
2020-10-05T10:57:19-0400
"v_{sound}=\sqrt{\frac{\gamma RT}{M}}\\v_{He}=\sqrt{\frac{5 RT}{3M}}=\sqrt{\frac{5 (8.31)(273)}{3(0.004)}}=972\frac{m}{s}"